OpenStudy (anonymous):

Match each equation with the ordered pair that is a solution. Match Term Definition y equals two x minus one A) one comma one y equals three x plus four B) three comma negative four y equals x plus five C) negative two comma three y equals x minus seven D) negative four comma negative eight

OpenStudy (anonymous):

1. y = 2x - 1 so (1,1) -- A 2. y = 3x + 4 so (-4,-8) -- D 3. y = x + 5 so (-2,3) -- C 4. y = x - 7 so (3,-4) -- B

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

You re welcome